The time, T, taken for a journey of a fixed distance is inversely proportional to the average speed S.


When S = S1, T = T1

On another journey, the speed S1 is increased by 40%


Work out the time taken for this journey.


Give your answer as a fraction of T1


1
Expert's answer
2021-02-01T08:22:51-0500

The time, T, taken for a journey of a fixed distance is inversely proportional to the average speed S.


"T_1=\dfrac{C}{S_1}"

On another journey, the speed S1 is increased by 40%


"S_2=1.4S_1"

Then


"T_2=\dfrac{C}{S_2}=\dfrac{C}{1.4S_1}=\dfrac{T_1}{1.4}"


"T_2=\dfrac{T_1}{1.4}"